# Rank1: Test-Time Compute for Reranking in Information Retrieval

### v1

@@ 5.2 Reasoning Language Models
Reasoning language models were introduced by OpenAI with their o1 model (Jaech et al., 2024). These models showed significantly improved performance on tasks that needed reasoning, such as math, logic, and programming. Since their release, many others have trained similar style models, including Google’s Gemini Flash Thinking and Deepseek’s R1. Notably, R1 is the only reasoning model that provides reasoning chains through APIs and is the only open-weights model.
@@
There have also been a flurry of works in the open-source space, both before and after o1 on reasoning language models and systems, focused on reproduction (Snell et al., 2024; Muennighoff et al., 2025), calibration and confidence (Jurayj et al., 2025), agentic capabilities with explainable reasoning traces (Weir et al., 2022, 2024), and much more. We expect this line of work to continue and continued collaboration will likely improve these models in retrieval as well.
@@ 6 Limitations and Future Work
Like other reasoning models, Rank1 can make mistakes and it can be surprising to see the model’s reasoning chain come close to a correct answer only to change it’s mind. We also found in particular that Rank1 can be particularly stringent in marking passages as true. For example, on a TREC COVID query about the origins of COVID-19 it marked every single passage as non-relevant, since none mentioned the specific wet market in Wuhan. However, when given a prompt to assume the user had no information about COVID-19, it was able to adapt better (although not perfectly). We observed this tendency to “overthink” when using the model interactively as it would already know the answer and was looking for a very specific phrase. We expect that this could be lessened with data that specifically trains the model to calibrate this.
@@
Rank1 also brings many more exciting areas for future work. As highlighted in the experiments section, it does extremely well despite the relatively small and non-diverse training data. Some of the straightforward areas of future work include:
@@
  • Listwise reasoning rerankers: although pointwise models are more efficient and more parallelizable, listwise rankers typically are more performant since they can see many documents at once. Simply gathering new data should enable this to be very successful.
@@ Appendix A Training and Hyperparameter Details
We fine-tune with LLaMA-Factory, using LoRA on all parameters with rank 32 and alpha 64. We use a learning rate of 1e-4 and an effective batch size of 128. We use early stopping based on the Bright Biology and NevIR scores. Models use prompts for BEIR and non-stackexchange cases so that they can understand the task (Appendix D).
@@
Appendix B Prompt for R1
@@
Prompt for Inference Determine if the following passage is relevant to the query. Answer only with ’true’ or ’false’. Query: query Passage: document ¡think¿
@@
Appendix C Examples of Unjudged and Incorrect DL19 Labels
@@
Passage: High blood pressure may also bring on heart failure by causing left ventricular hypertrophy, a thickening of the heart muscle that results in less effective muscle relaxation between heart beats. This makes it difficult for the heart to fill with enough blood to supply the body’s organs, especially during exercise, leading your body to hold onto fluids and your heart rate to increase.
@@
Appendix D Prompts for specific datasets
@@
SciFact Claim: FILL_QUERY_HERE¡newline¿¡newline¿A relevant passage would provide evidence that either **supports** or **refutes** this claim. A passage with any information on any related subpart should be relevant.
ClimateFEVER Claim: FILL_QUERY_HERE¡newline¿¡newline¿A relevant passage would provide evidence that either **supports** or **refutes** this claim. A passage with any information on any related subpart should be relevant.
@@
ArguAna I am looking to write an essay and need to find counterarguments against this statement:¡newline¿¡newline¿FILL_QUERY_HERE¡newline¿¡newline¿Does this passage have any counterargument or evidence that could be used to help me?
DBPedia I am looking to write an essay on this topic and need as much related background information to help me. The topic is:¡newline¿¡newline¿FILL_QUERY_HERE¡newline¿¡newline¿If the passage provides any background information that could be connected it is relevant.
@@
NFCorpus Topic: FILL_QUERY_HERE¡newline¿¡newline¿Given the above topic, I need to learn about all aspects of it. It does not need to be directly relevant, only tangentially informational. Please mark as relevant any passages with even weak connections. I need to learn fast for my job, which means I need to understand each part individually.¡newline¿¡newline¿Again remember, any connection means relevant even if indirect. So if it is not addressed, that is okay – it does not need to be explicitly.¡newline¿¡newline¿Find me passages with any type of connection, including weak connections!!!!
@@
BrightRetrieval aops Find different but similar math problems to FILL_QUERY_HERE¡newline¿¡newline¿A document is relevant if it uses the same class of functions and shares **any** overlapping techniques.
@@
BrightRetrieval leetcode I am looking to find different problems that share similar data structures (of any kind) or algorithms (e.g. DFS, DP, sorting, traversals, etc.). I am looking for problems that share one or both of these similarities to this:¡newline¿¡newline¿FILL_QUERY_HERE¡newline¿¡newline¿Does this passage share any similarities? e.g. if there was a textbook on leetcode problems, this would be in the same book even though it could be in a different chapter.
BrightRetrieval pony I will use the programming language pony. Problem: FILL_QUERY_HERE¡newline¿¡newline¿But to solve the problem above, I need to know things about pony. A passage is relevant if it contains docs that match any part (even basic parts) of the code I will have to write for the above program.
BrightRetrieval Can you find background information about the concepts used to answer the question:¡newline¿¡newline¿FILL_QUERY_HERE¡newline¿¡newline¿A passage is relevant if it contains background information about a **sub-concept** that someone might cite/link to when answering the above question.
@@
Appendix E Things we tried that didn’t work
We tried a few other things that didn’t work:
  • We tried to calibrate the scores better by adding a ModernBERT model (Warner et al., 2024) on top of the outputs of the reasoning chain. However, this ended up performing worse. It is possible that with better curation this would be an effective approach however.
  • We tried adding extra loss to the last token of the next token prediction loss in LLaMA-Factory. However, this resulted in sub-par performance, likely because predicting the last true/false token at the end of reasoning chain is fairly easy (as the model already states beforehand “the answer is X” leaving little room for doubt).
Appendix F Noise in BEIR benchmarks
@@
In the original SciFact work (Wadden et al., 2020) the authors have three labels: support, refute, or not enough information (NEI). However, when incorporated into BEIR, the NEI queries are still included. As there was not evidence to support these queries (as determined by the original authors) these queries are effectively noise. In practice, what it means is that tests model’s ability to find the top ranked BM25 document, which is guaranteed to not have enough information to either refute or support the claim. The large number of these queries add a significant amount of noise.
@@
FiQA is scraped from Financial Stackexchange (Maia et al., 2018). However, during the original scrape the creators did not collect the post’s details. Thus, the retrieval setup is to take the posts title and search for the top answer. Yet, in many cases the user clarified important details in the post that entirely changed the meaning of the query. Without this additional information for some queries it is impossible for someone to determine what the best answer is as the post details asked many other questions that were different from the title of the post.
@@
DBPedia has many partially relevant labels, that give credit for finding non-relevant information. For example, the entire passage with a relevance of “1” for the query ”Eiffel Tower” is “The year 1989 in architecture involved some significant architectural events and new buildings.” However, nothing about the Eiffel Tower is connected to 1989. After digging, the only connection is its the 100 year anniversary of it being built – but the passage does not mention this. There are many such examples in the dataset (and more documents judged “1” than “2”), contributing to the noise.
@@
Touche2020 has been well examined by Thakur et al. (2024). They created a much cleaner version of the data, however, it is not the “standard” evaluation set in BEIR.
@@
Many of the other datasets give credit for partial relevance in a similar manner to DBPedia. This includes TREC COVID (which has so many real positives that it is a non-issue for the top-10), NFCorpus, and SciDocs. For many of these datasets, it is very difficult for even a human to match relevance: e.g. on NFCorpus you would have to guess any potential link that had been on that website and was even ancillarily related to the title. An example is the query “How Fruits and Vegetables Can Treat Asthma” which matches a document with the title “Effect of a single high-fat meal on endothelial function in healthy subjects.” and does not discuss asthma at all. The reranker would have to assume that any passage discussing food or asthma separately in any context would be relevant. Although one intution is that these documents should rank higher than completely non-relevant documents, for Rank1 it treats them as the same (as they are both equally non-relevant to the query). This may be suboptimal for some approaches, but for today’s RAG use cases returning only actual relevant documents seems more useful.

### v2

@@ 5.2 Reasoning Language Models
Reasoning language models were introduced by OpenAI with their o1 model (Jaech et al., 2024). These models showed significantly improved performance on tasks that needed reasoning, such as math, logic, and programming. Since their release, many others have trained similar style models, including Google's Gemini Flash Thinking and Deepseek's R1. Notably, R1 is the only reasoning model that provides reasoning chains through APIs and is the only open-weights model.
@@
There have also been a flurry of works in the open-source space, both before and after o1 on reasoning language models and systems, focused on reproduction (Snell et al., 2024; Muennighoff et al., 2025), calibration and confidence (Jurayj et al., 2025), agentic capabilities with explainable reasoning traces (Weir et al., 2022; 2024), and much more. We expect this line of work to continue and continued collaboration will likely improve these models in retrieval as well.
@@ 6 Limitations and Future Work
Like other reasoning models, Rank1 can make mistakes and it can be surprising to see the model's reasoning chain come close to a correct answer only to change its mind. We also found that Rank1 can be particularly stringent in marking passages as true. For example, on a TREC COVID query about the origins of COVID-19 it marked every single passage as non-relevant, since none mentioned the specific wet market in Wuhan. However, when given a prompt to assume that the user had no information about COVID-19, it was able to adapt better (although not perfectly). We observed this tendency to ``overthink" when using the model interactively, as it would already know the answer and was looking for a very specific phrase. We expect that this could be reduced with data that specifically trains the model to calibrate this.
@@
Rank1 also brings many more exciting areas for future work (and concurrent work (Yang et al., 2025)). As highlighted in the experiments section, it does extremely well despite the relatively small and non-diverse training data. Some of the promising areas of future work are:
@@
  • Listwise reasoning rerankers: although pointwise models are more efficient and more parallelizable, listwise rerankers are generally more performant since they can see many documents at once. Simply gathering new data should enable this to be very successful.
@@
We expect to see many future applications of test-time compute applied to retrieval with great success, as the recipe is both simple and effective.
@@ 7 Conclusion
8 Ethics Statement
Rank1 is a more generative form of many rerankers currently used in information retrieval. As such, it introduces the possibility of generating flawed text (just like LMs). This includes incorrect and hallucinated information, the potential for toxic language, biased output, and other such flaws. These same flaws may also be present in the R1-generated dataset used to train these models.
As an improved reranker, is also increases the risk when used for illicit activities. As shown by (BehnamGhader et al., 2025) models that are better at ranking and following instructions can be used more effectively to retrieve information that leads to illicit purposes. However, most of the use cases of information retrieval are positive; such is the balance when using a dual-use technology. We would encourage users to use improved retrieval capabilities for positive purposes.
@@ Appendix A Training and Hyperparameter Details
We fine-tune with LLaMA-Factory, using LoRA on all parameters with rank 32 and alpha 64. We use a learning rate of 1e-4 and an effective batch size of 128. We use early stopping based on the Bright Biology and NevIR scores. Models use prompts for BEIR and non-stackexchange cases so that they can understand the task (Appendix E).
@@
Appendix B Data Analysis from mT5 Negatives
We sample 20 of the instances from mT5 that were ``hard negatives" and manually evaluated them. We found that 12 of them were in fact positives and thus false negatives. This helps to explain why filtering these out made such a large difference in the training data, as roughly 2/3rds of these mT5 negatives were actually positive, making the training process more noisy.
Appendix C Prompt for R1
@@
Prompt for Inference Determine if the following passage is relevant to the query. Answer only with 'true' or 'false'. Query: query Passage: document <think>
@@
Appendix D Examples of Unjudged and Incorrect DL19 Labels
@@
Passage: High blood pressure may also bring on heart failure by causing left ventricular hypertrophy, a thickening of the heart muscle that results in less effective muscle relaxation between heart beats. This makes it difficult for the heart to fill with enough blood to supply the body's organs, especially during exercise, leading your body to hold onto fluids and your heart rate to increase.
@@
Appendix E Prompts for specific datasets
@@
SciFact Claim: FILL_QUERY_HERE<newline><newline>A relevant passage would provide evidence that either **supports** or **refutes** this claim. A passage with any information on any related subpart should be relevant.
ClimateFEVER Claim: FILL_QUERY_HERE<newline><newline>A relevant passage would provide evidence that either **supports** or **refutes** this claim. A passage with any information on any related subpart should be relevant.
@@
ArguAna I am looking to write an essay and need to find counterarguments against this statement:<newline><newline>FILL_QUERY_HERE<newline><newline>Does this passage have any counterargument or evidence that could be used to help me?
DBPedia I am looking to write an essay on this topic and need as much related background information to help me. The topic is:<newline><newline>FILL_QUERY_HERE<newline><newline>If the passage provides any background information that could be connected it is relevant.
@@
NFCorpus Topic: FILL_QUERY_HERE<newline><newline>Given the above topic, I need to learn about all aspects of it. It does not need to be directly relevant, only tangentially informational. Please mark as relevant any passages with even weak connections. I need to learn fast for my job, which means I need to understand each part individually.<newline><newline>Again remember, any connection means relevant even if indirect. So if it is not addressed, that is okay – it does not need to be explicitly.<newline><newline>Find me passages with any type of connection, including weak connections!!!!
@@
BrightRetrieval aops Find different but similar math problems to FILL_QUERY_HERE<newline><newline>A document is relevant if it uses the same class of functions and shares **any** overlapping techniques.
@@
BrightRetrieval leetcode I am looking to find different problems that share similar data structures (of any kind) or algorithms (e.g. DFS, DP, sorting, traversals, etc.). I am looking for problems that share one or both of these similarities to this:<newline><newline>FILL_QUERY_HERE<newline><newline>Does this passage share any similarities? e.g. if there was a textbook on leetcode problems, this would be in the same book even though it could be in a different chapter.
BrightRetrieval pony I will use the programming language pony. Problem: FILL_QUERY_HERE<newline><newline>But to solve the problem above, I need to know things about pony. A passage is relevant if it contains docs that match any part (even basic parts) of the code I will have to write for the above program.
BrightRetrieval Can you find background information about the concepts used to answer the question:<newline><newline>FILL_QUERY_HERE<newline><newline>A passage is relevant if it contains background information about a **sub-concept** that someone might cite/link to when answering the above question.
@@
Appendix F Things we tried that didn't work
We tried a few other things that didn't work:
  • We tried to calibrate the scores better by adding a ModernBERT model (Warner et al., 2024; Weller et al., 2025b) on top of the outputs of the reasoning chain. However, this ended up performing worse. It is possible that with better curation this would be an effective approach however.
  • We tried adding extra loss to the last token of the next token prediction loss in LLaMA-Factory. However, this resulted in sub-par performance, likely because predicting the last true/false token at the end of reasoning chain is fairly easy (as the model already states beforehand ``the answer is X" leaving little room for doubt).
Appendix G Noise in BEIR benchmarks
@@
In the original SciFact work (Wadden et al., 2020) the authors have three labels: support, refute, or not enough information (NEI). However, when incorporated into BEIR, the NEI queries are still included. As there was not evidence to support these queries (as determined by the original authors) these queries are effectively noise. In practice, what it means is that tests model's ability to find the top ranked BM25 document, which is guaranteed to not have enough information to either refute or support the claim. The large number of these queries add a significant amount of noise.
@@
FiQA is scraped from Financial Stackexchange (Maia et al., 2018). However, during the original scrape the creators did not collect the post's details. Thus, the retrieval setup is to take the posts title and search for the top answer. Yet, in many cases the user clarified important details in the post that entirely changed the meaning of the query. Without this additional information for some queries it is impossible for someone to determine what the best answer is as the post details asked many other questions that were different from the title of the post.
@@
DBPedia has many partially relevant labels, that give credit for finding non-relevant information. For example, the entire passage with a relevance of ``1" for the query "Eiffel Tower" is ``The year 1989 in architecture involved some significant architectural events and new buildings." However, nothing about the Eiffel Tower is connected to 1989. After digging, the only connection is its the 100 year anniversary of it being built – but the passage does not mention this. There are many such examples in the dataset (and more documents judged ``1" than ``2"), contributing to the noise.
@@
Touche2020 has been well examined by Thakur et al. (2024). They created a much cleaner version of the data, however, it is not the ``standard" evaluation set in BEIR.
@@
Many of the other datasets give credit for partial relevance in a similar manner to DBPedia. This includes TREC COVID (which has so many real positives that it is a non-issue for the top-10), NFCorpus, and SciDocs. For many of these datasets, it is very difficult for even a human to match relevance: e.g. on NFCorpus you would have to guess any potential link that had been on that website and was even ancillary related to the title. An example is the query ``How Fruits and Vegetables Can Treat Asthma" which matches a document with the title ``Effect of a single high-fat meal on endothelial function in healthy subjects." and does not discuss asthma at all. The reranker would have to assume that any passage discussing food or asthma separately in any context would be relevant. Although one intuition is that these documents should rank higher than completely non-relevant documents, for Rank1 it treats them as the same (as they are both equally non-relevant to the query). This may be suboptimal for some approaches, but for today's RAG use cases returning only actual relevant documents seems more useful.